OpenStudy (anonymous):

Find the indefinite integral: Integrand of (4x^3+2)/((x^4+2x)^2) dx

OpenStudy (anonymous):

u=(x^4+2x), du= 12x^2 dx

OpenStudy (anonymous):

how do I simplify from here since its a fraction?

OpenStudy (anonymous):

that'll do it!

OpenStudy (anonymous):

well actually it is \(u=x^4+2x,du=(4x^3+2)dx\)

OpenStudy (anonymous):

then integral becomes \[\int \frac{du}{u^2}\] use the power rule backwards to solve

OpenStudy (anonymous):

umm.. wouldn't du be 12x^2 since that is the derivative of (4x^3+2)?

OpenStudy (anonymous):

so I would do the anti derivative of (du)/u^2

OpenStudy (anonymous):

wait wait you do not take the derivative twice, just once

OpenStudy (anonymous):

\[u(x)=x^4+2x\] \[\frac{du}{dx}=4x^3+2\] \[du=(4x^3+2)dx\] like that, which as luck would have it is exactly what you have in your problem

OpenStudy (anonymous):

so yes, you take the anti derivative of \(\frac{1}{u^2}\) which is straight forward

OpenStudy (anonymous):

i see..

OpenStudy (anonymous):

you get \(-\frac{1}{u}\) by the power rule backwards, then replace \(u\) by \(x^4+2x\) for your answer

OpenStudy (anonymous):

would my answer be -(x^4+2x)^(-1)+C?

OpenStudy (anonymous):

-1/(x^4+2 x)+constant

OpenStudy (anonymous):

oh ok.. i move (x^4+2x) down to be the denominator.

OpenStudy (anonymous):

yes that is one way to write it with exponential notation easier way to understand is \[-\frac{1}{x^4+2x}+C\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

one question which gets me confused. does the "du" get dropped when I do the anti derivative?

OpenStudy (anonymous):

I always think i need to include it in the final simplified answer

OpenStudy (anonymous):

what you should do to understand this trick of notation is take the derivative of your answer, and see why it gives you exactly what you want it is the chain rule backwards

OpenStudy (anonymous):

yes, the du gets dropped, it just says what the variable is, like the notation \[\frac{d}{dx}[ax^2]\] for example

OpenStudy (anonymous):

thank you